OpenStudy (anonymous):

find the simple interest and total amounts paid on the loan principal$450.0 Rate 5% Time 3yrs what is the interest and amount please explain in simplest terms

OpenStudy (callisto):

Interest of the loan = 450 (5%) x3 = 22.5 x3 = 67.5 Total amount paid = interest + principle = 67.5 + 450 = work it out yourself :)

OpenStudy (anonymous):

find the annual rate of interest from principal $1500.0 interest $210.0 1 yr what is the rate

OpenStudy (callisto):

let r be the interest rate 1500 (r) = 210 solve it and convert r into percentage

OpenStudy (callisto):

actually, you should know what principle * interest rate = interest

OpenStudy (anonymous):

is it 31.50

OpenStudy (anonymous):

dont get it?

OpenStudy (anonymous):

$1,500(r)=$210 Divide by sides by $1,500 So you have $1,500/$1,500 (r)=$210/$1,500 which becomes 1(r)=.14 so the rate is r=.14 To convert a decimal to a percent, move the decimal place to the right two places and attach a percent sign. So .14 becomes 14. Now attach the percent sign. So 14. becomes 14.% or 14%

OpenStudy (callisto):

r = 210/1500 *100% = 0.14 *100% = 14%